OpenStudy (anonymous):

find three consecutive even integers such that the first times the third is 2 greater the 5 times the second?

terenzreignz (terenzreignz):

http://openstudy.com/study#/updates/4feae23ae4b0bbec5cfaa74d

OpenStudy (radar):

Let 2x = first even integer then 2x + 2 = 2nd even integer and 2x+4 =3rd even integer Statement says: 2x(2x+4)=2+5(2x+2) \[4x ^{2}+8x=2+10x+10\]\[4x ^{2}-2x-12=0\] \[2x ^{2}-x-6=0\]\[(x-2) (2x+3)=0\] x=2 2x=4 (first positive digit) 2x+2= 6 (second positive digit) 2x+4= 8 third positiv digit Check:4*3=2+(5*2) 12=12 Now for the other root where x=-3/2 is 2x=-3 is not even so disregard.